OpenStudy (anonymous):

Solve the equation using square roots. 5x^2-45=0

OpenStudy (anonymous):

\[5x^2=45\\ x^2=\frac{45}{5}\\ x=\pm \sqrt{ 9} \\ x= \pm 3\]